# On the effects of radiation on mass transfer in binary stars

### v1

@@ 4.5 Comparison to previous works
We now compare the approach in this paper to our previous work in CP23. For convenience, we summarize the equations from CP23 in Appendix C. The greatest difference between these works comes from the fact that here we assume that all physical quantities are constant along $Q$ and we use a physically motivated expression for $Q$ (equations 30, 31), where $Q$ is appropriately enlarged by a factor of $1/(1-\Gamma_{\rm Edd})$. In contrast, in CP23 we assumed hydrostatic equilibrium in the $yz$-plane together with a polytropic EOS and averaged over the $yz$-plane. As a result, characteristic cross sections corresponding to density and pressure profiles, $Q_{\rho}$ and $Q_{P}$, appeared in all equations. Moreover, CP23 had an extra term on the right-hand side of the energy equation (63c) that is a consequence of the average of the Roche potential $\phi_{\rm R}$ in the $yz$ plane. The effects of this term can be understood by realizing that if the gas is farther from the $x$ axis, its potential is greater. Consequently, if the cross section shrinks, the potential energy becomes lower, and this energy difference is used to heat the gas. Thus, our original CP23 model is not strictly adiabatic but is adiabatic in the sense that there is no energy exchange between the gas and radiation.
In the present paper, we define two limiting cases: the small-$v$ case that is inspired by the hydrostatic case (Section 3.2) and the adiabatic case with no energy exchange between gas and radiation and no extra term on the right-hand side of the energy equation (Section 3.3). The adiabatic case behaves similarly to the original “adiabatic” CP23 model with the difference that the adiabatic case gives a steeper temperature gradient resulting in a lower temperature at L1 and thus lower critical velocity $v_{\rm crit}$ and a lower value of the MT rate. Also, if $x_{0}$ approaches $x_{1}$, the two cases obtain different $Q$ and $v_{\text{crit}}$ and consequently also different MT rates. Specifically, the adiabatic case obtains $Q\propto P_{\rm gas}/\rho/(1-\Gamma_{\rm Edd})\propto P/\rho$ and $v_{\rm crit}^{2}=P_{\rm gas}/\rho$, while our original model had $Q\propto c_{T,\rm gas}^{2}\propto P_{\rm gas}/\rho$ and $v_{\rm crit}^{2}\sim P/\rho$. Hence, the original model obtains a lower value of the MT rate even in this limiting case as $-\dot{M}_{\rm d}\propto v_{\rm crit}Q$. The small-$v$ case and the adiabatic case obtain the same MT rate in the limit of $x_{0}\to x_{1}$ because they predict the same $Q$ and $v_{\text{crit}}$. This behavior is illustrated in Figs. 2 and 3.
@@ 5 Super-Eddington mass transfer
We now analyze situations where the star exceeds the Eddington limit interior of the L1 point. We could still solve our equations with $x_{0}$ in a sub-Eddington region and define $x_{1}$ as the point where $\Gamma_{\text{Edd}}=1$. However, super-Eddington flux can often be carried out by convection without any outflows. Outflows may only develop when convective efficiency diminishes and $v_{\text{conv}}$ approaches the sound speed. In principle, our equations could be extended to include convective energy transport using, for example, the formalism of Kuhfuss (1986), where the outflow should naturally develop once convection is unable to carry the flux. However, studies of single nonrotating super-Eddington stars have shown that multidimensional effects fundamentally alter the structure of the relevant regions (e.g., Shaviv, 2001b; Owocki et al., 2004; Jiang et al., 2015, 2018). This suggests that further refinements of our 1D model may be of limited value. Instead, we revisit some of the arguments constraining super-Eddington outflows in spherical symmetry and find that proximity to the L1 point may relax some of them. While our results are promising with respect to the potential for a super-Eddington MT boost, determining whether such outflows actually occur will ultimately require multi-dimensional time-dependent simulations.
@@ 5.1 Analytic theory
For the discussion of the two remaining issues, we first outline the theory of super-Eddington outflows from spherically-symmetric stars presented by Owocki et al. (2004) and later revisited by Quataert et al. (2016) and Owocki et al. (2017), and then modify the results to the context of the L1 point. For $\Gamma_{\text{tot}}=L/L_{\text{Edd}}>1$ and inefficient convection, the entire luminosity can be carried by an outflow with velocity $v_{\text{out}}\gtrsim c_{T,\text{gas}}$ satisfying
@@
where $\alpha_{\text{tir}}=1-1/\Gamma_{\text{tot}}$ is the efficiency parameter of photon tiring (Owocki et al., 2017). If $\dot{M}_{\text{out}}\gtrsim\dot{M}_{\text{tir}}$, the outflow should just stall. The condition for an outflow to form is thus
@@
Figure 6: Radial profiles of key quantities for a $30\,\rm{M}_{\odot}$ low-metallicity star during the super-Eddington MT boost at $t=6.972$ Myr. The upper left panel shows Eddington ratios of total ($L$) and radiative ($L_{\text{rad}}$) luminosities and density and temperature profiles. Bottom left panel shows values of escape velocity $v_{\text{esc}}$, sound speeds $c_{S}$ and $c_{T,\text{gas}}$, convective velocity reported by the MLT modules in MESA $v_{\rm conv}$, outflow velocity $v_{\text{out}}$ (equation 42), and the critical velocity $v_{\text{crit}}$ of Quataert et al. (2016). Top right panel shows estimates of L1 MT rates based on super-Eddington boosted MT $\dot{M}_{\text{out,L1}}$ (equation 46), a more pessimistic estimate with $v_{\text{out}}$ replaced with $c_{T,\text{gas}}$, $L_{\rm d}/v_{\rm esc}^{2}$, and the prediction of Marchant et al. (2021) $\dot{M}_{\text{M}}$. In addition, we show also the photon tiring limit $\dot{M}_{\text{tir,L1}}$ (equation 47). Bottom right panel displays the estimate for L1 nozzle area $c_{T,\text{gas}}^{2}/v_{\text{esc}}^{2}$, relative distance to the L1 point $\delta R_{\text{L}}$, efficiency parameter of photon tiring $\alpha_{\text{tir}}$, and the condition for super-Eddington boost $\beta_{\text{SE,L1}}$ evaluated using $v_{\text{out}}$ or $c_{T,\text{gas}}$ (equation 48). In all panels, light and dark shading marks super-Eddington regions based on $L$ and $L_{\text{rad}}$, vertical dashed lines shows the interior most point where $v_{\text{conv}}>c_{S}$, and the vertical dotted line indicates the minimum of $\beta_{\text{SE, L1}}$.
Models based on the porosity length formalism suggest that the maximum achievable mass loss rate is proportional to the gravitational scale height (e.g. Owocki et al., 2004), and thus inversely proportional to effective gravity $g_{\rm eff}\propto\nabla\phi_{\rm R}$ in a Roche model, where $\nabla\phi_{\rm R}\to 0$ close to L1. Consequently, we assume that if an outflow occurs from a star nearly filling its Roche lobe and possessing a super-Eddington, convectively inefficient subsurface zone, it will primarily flow through the nozzle surrounding L1 at a rate $\dot{M}_{\text{out,L1}}$. The detailed steps needed for the modifications from spherical symmetry to the vicinity of L1 are described in Appendix E. In summary, we take into account the limited extent of the nozzle by multiplying $\dot{M}$ by $Q/(4\pi R_{\text{d}}^{2})\sim c_{T,\text{gas}}^{2}/v_{\text{esc}}^{2}$, and differences in potential proportional to $v_{\text{esc}}^{2}$ are multiplied by $\delta R_{\text{L}}$. These modifications give
@@
Since $\delta R_{\text{L}}$ can be made arbitrarily small by positioning the super-Eddington layer very close to the Roche lobe, this modified condition can be satisfied and the super-Eddington outflow will not be limited by photon tiring. In such a case, a super-Eddington outflow with velocity $v_{\rm out}$ and MT rate $-\dot{M}_{\rm d}\sim\dot{M}_{\rm out,L1}$ forms through the L1 point.
Quataert et al. (2016) presented a different formulation of the photon tiring limit. They argued that equation (42) is valid when the critical velocity of the outflow, $v_{\text{crit}}\sim(GL_{\text{d}})^{1/5}$, exceeds the local escape velocity modified by a factor taking into account the structure of the surrounding envelope. In the vicinity of the L1 point, their condition reads $v_{\text{crit}}\gtrsim v_{\text{esc}}\delta R_{\text{L}}^{1/2}$. We show this condition in the figures that illustrate the super-Eddington MT boost, but we do not find significant differences in the outcomes when compared to equation (48).
@@ 5.2 Application to a realistic star
To illustrate the super-Eddington MT boost we use the same $30\,\rm{M}_{\odot}$ low-metallicity donor as in Sec. 4.4.1 but this time in a binary with an initial $4.5\,\rm{M}_{\odot}$ black hole and initial orbital period of $1000$ days. Our first goal is to investigate how to evaluate the conditions for a super-Eddington MT boost on realistic stellar profiles. In Fig. 6, we show radial profiles of a number of relevant quantities near the outer edge of the stellar envelope at a moment, when the donor still underfills its Roche lobe. We see that the super-Eddington layer reaches $\Gamma_{\text{tot}}\gtrsim 10$, occurs in the hydrogen recombination zone, and is accompanied by a density inversion (upper left panel). We find that simple estimates of the convective energy flux differ sufficiently from the results of mixing length theory (MLT) of convection implemented in MESA and are not useful. Instead, we diagnose the inefficiency of convection by comparing the $v_{\text{conv}}$ reported by MLT with the local sound speed (lower left panel), finding that convection is inefficient in a region $0.990\lesssim r/R_{\text{L}}\lesssim 0.993$. This result is insensitive to the use of $c_{S}$ or $c_{T,\text{gas}}$ for comparison purposes. The lower left panel of Fig. 6 also shows $v_{\text{crit}}$. Near the L1 point, we find that the relevant quantities for the Quataert et al. (2016) version of the photon tiring are $v_{\text{crit}}\approx 25\,$km s-1, $v_{\text{esc}}\approx 120\,$km s-1, and $\delta R_{\text{L}}^{1/2}\approx 0.1$ implying that the requirement is satisfied and equation (42) can be used to find $v_{\text{out}}$. Our results indicate that $v_{\text{out}}$ exceeds the local sound speed by a factor of 3–4, but remains smaller than $v_{\text{esc}}$. The upper right panel of Fig. 6 displays $\dot{M}$ near L1 corresponding to the MT flow and the photon tiring limit. The shape of $\dot{M}_{\text{tir}}$ is mostly determined by $\alpha_{\text{tir}}$ (bottom right panel). We find that $\dot{M}_{\text{out,L1}}<\dot{M}_{\text{tir,L1}}$ and correspondingly $\beta_{\text{SE,L1}}<1$ in a significant range of $r$. Since there is some ambiguity about the correct value of $v_{\text{out}}$, we also show an estimate of $\dot{M}=(c_{T,\text{gas}}/v_{\text{esc}})^{2}L_{\rm d}/c_{T,\rm gas}^{2}=L_{\rm d}/v_{\text{esc}}^{2}$ and modified $\beta_{\text{SE,L1}}(c_{T,\text{gas}})$, where $v_{\text{out}}$ is replaced with $c_{T,\text{gas}}$ in equation (48). Even in this more pessimistic case, the implied MT rate does not exceed the photon tiring limit. While in this case the outflow condition is satisfied later in the evolution (only about 700 years before Roche-lobe overflow), the implied MT rates are higher leading to a higher total mass lost.
@@
Figure 7: Time evolution of donor properties illustrating super-Eddington boosted MT on an initially $30\,\rm{M}_{\odot}$ star in a binary with a $4.5\,\rm{M}_{\odot}$ black hole. Top panel shows estimates of L1 MT rates $\dot{M}_{\text{out,L1}}$ (equation 46), $\dot{M}_{\text{M}}$ (Marchant et al., 2021), and a more pessimistic estimate for super-Eddington MT $\dot{M}=L_{\rm d}/v_{\text{esc}}^{2}$, as well as photon tiring limit $\dot{M}_{\text{tir,L1}}$ and wind mass loss from the donor $\dot{M}_{\text{wind}}$. Second panel shows the cumulative amounts of mass lost from the donor for MT rates shown in the top panel. Third panel shows the evolution of relevant velocities $v_{\text{out}}$, $v_{\text{esc}}$, $v_{\text{crit}}$, and $c_{S}$. Bottom panel shows time evolution of the minimum of $\beta_{\text{SE,L1}}$, maximum of $v_{\text{conv}}/c_{S}$, and maximum of $\Gamma_{\text{tot}}=L/L_{\rm Edd}$ evaluated over the subsurface region that could drive super-Eddington MT, and the relative Roche-lobe-filling factor of the star $1-R_{\text{d}}/R_{\text{L}}$.
@@
To summarize Fig. 6, we find that at this point in evolution the donor is underfilling its Roche lobe, but the super-Eddington flux in the subsurface layer probably leads to MT through L1 with a rate of around $10^{-2}\,\rm{M}_{\odot}\,\text{yr}^{-1}$, which is significantly lower than the photon tiring limit near L1 but also significantly higher than the MT rate of about $10^{-4}\,\rm{M}_{\odot}\,\text{yr}^{-1}$ calculated using the traditional approach (Marchant et al., 2021).
[…]
@@
By integrating $\dot{M}_{\text{d}}$, we find that the total mass lost from the donor due to the super-Eddington MT exceeds $10\,\rm{M}_{\odot}$. We obtain even higher numbers for the more pessimistic case when $v_{\text{out}}=c_{T,\text{gas}}$ or similar numbers even when we restrict the super-Eddington MT to the final $1$ kyr before Roche-lobe overflow when $v_{\text{crit}}>v_{\text{esc}}\delta R_{\text{L}}^{1/2}$. These numbers are clearly an artifact of not evolving the binary self-consistently, however, they suggest that the super-Eddington MT boost can easily overwhelm traditional mode of MT and can cause significant deviations in binary evolution pathways.

### v2

@@ 4.5 Comparison to previous works
We now compare the approach in this paper to our previous work in cehula2023. For convenience, we summarize the equations from cehula2023 in Appendix C. The greatest difference between these works comes from the fact that here we assume that all physical quantities are constant along $Q$ and we use a physically motivated expression for $Q$ (equations 30, 31), where $Q$ is appropriately enlarged by a factor of $1/(1-\Gamma_{\rm Edd})$. In contrast, in cehula2023 we assumed hydrostatic equilibrium in the $yz$-plane together with a polytropic EOS and averaged over the $yz$-plane. As a result, characteristic cross sections corresponding to density and pressure profiles, $Q_{\rho}$ and $Q_{P}$, appeared in all equations. Moreover, cehula2023 had an extra term on the right-hand side of the energy equation (63c) that is a consequence of the average of the Roche potential $\phi_{\rm R}$ in the $yz$ plane. The effects of this term can be understood by realizing that if the gas is farther from the $x$ axis, its potential is greater. Consequently, if the cross section shrinks, the potential energy becomes lower, and this energy difference is used to heat the gas. Thus, our original cehula2023 model is not strictly adiabatic but is adiabatic in the sense that there is no energy exchange between the gas and radiation.
In the present paper, we define two limiting cases: the small-$v$ case that is inspired by the hydrostatic case (Section 3.2) and the adiabatic case with no energy exchange between gas and radiation and no extra term on the right-hand side of the energy equation (Section 3.3). The adiabatic case behaves similarly to the original “adiabatic” cehula2023 model with the difference that the adiabatic case gives a steeper temperature gradient resulting in a lower temperature at L1 and thus lower critical velocity $v_{\rm crit}$ and a lower value of the MT rate. Also, if $x_{0}$ approaches $x_{1}$, the two cases obtain different $Q$ and $v_{\text{crit}}$ and consequently also different MT rates. Specifically, the adiabatic case obtains $Q\propto P_{\rm gas}/\rho/(1-\Gamma_{\rm Edd})\propto P/\rho$ and $v_{\rm crit}^{2}=P_{\rm gas}/\rho$, while our original model had $Q\propto c_{T,\rm gas}^{2}\propto P_{\rm gas}/\rho$ and $v_{\rm crit}^{2}\sim P/\rho$. Hence, the original model obtains a lower value of the MT rate even in this limiting case as $\dot{M}_{\rm d}\propto v_{\rm crit}Q$. The small-$v$ case and the adiabatic case obtain the same MT rate in the limit of $x_{0}\to x_{1}$ because they predict the same $Q$ and $v_{\text{crit}}$. This behavior is illustrated in Figs. 2 and 3.
@@ 5 Super-Eddington mass transfer
We now analyze situations where the star exceeds the Eddington limit interior of the L1 point. We could still solve our equations with $x_{0}$ in a sub-Eddington region and define $x_{1}$ as the point where $\Gamma_{\text{Edd}}=1$. However, super-Eddington flux can often be carried out by convection without any outflows. Outflows may only develop when convective efficiency diminishes and $v_{\text{conv}}$ approaches the sound speed. In principle, our equations could be extended to include convective energy transport using, for example, the formalism of kuhfuss1986, where the outflow should naturally develop once convection is unable to carry the flux. However, studies of single nonrotating super-Eddington stars have shown that multidimensional effects fundamentally alter the structure of the relevant regions (e.g., shaviv01a; owocki2004; jiang2015; jiang2018). This suggests that further refinements of our 1D model may be of limited value. Instead, we revisit some of the arguments constraining super-Eddington outflows in spherical symmetry and find that proximity to the L1 point may relax some of them. While our results are promising with respect to the potential for a super-Eddington MT boost, determining whether such outflows actually occur will ultimately require multi-dimensional time-dependent simulations.
@@ 5.1 Analytic theory
For the discussion of the two remaining issues, we first outline the theory of super-Eddington outflows from spherically-symmetric stars presented by owocki2004 and later revisited by quataert16 and owocki2017, and then modify the results to the context of the L1 point. For $\Gamma_{\text{tot}}=L/L_{\text{Edd}}>1$ and inefficient convection, the entire luminosity can be carried by an outflow with velocity $v_{\text{out}}\gtrsim c_{T,\text{gas}}$ satisfying
@@
where $\alpha_{\text{tir}}=1-1/\Gamma_{\text{tot}}$ is the efficiency parameter of photon tiring (owocki2017). If $\dot{M}_{\text{out}}\gtrsim\dot{M}_{\text{tir}}$, the outflow should just stall. The condition for an outflow to form is thus
@@
Figure 6: Radial profiles of key quantities for an initially $30\,\rm{M}_{\odot}$ low-metallicity star during the super-Eddington MT boost at $t=6.972$ Myr, when the donor has $M_{\text{d}}=28.57\,\rm{M}_{\odot}$, $T_{\text{eff}}=5255$ K, $L_{\text{d}}=3.96\times 10^{5}\,\rm{L}_{\odot}$, and $R_{\text{d}}=759\,\rm{R}_{\odot}$. Upper left panel shows profiles of density, temperature, and opacity. Middle left panel shows Eddington ratios of total ($L$) and radiative ($L_{\text{rad}}$) luminosities and relative luminosity profile $l/L_{\text{d}}$. Bottom left panel shows values of escape velocity $v_{\text{esc}}$, sound speeds $c_{S}$ and $c_{T,\text{gas}}$, convective velocity reported by the MLT modules in MESA $v_{\rm conv}$, outflow velocity $v_{\text{out}}$ (equation 42), and the critical velocity $v_{\text{crit}}$ of quataert16. Upper right panel displays the estimate for L1 nozzle area $c_{T,\text{gas}}^{2}/v_{\text{esc}}^{2}$, relative distance to the L1 point $\delta R_{\text{L}}$, and efficiency parameter of photon tiring $\alpha_{\text{tir}}$. Middle right panel shows estimates of L1 MT rates based on super-Eddington boosted MT $\dot{M}_{\text{out,L1}}$ (equation 46), a more pessimistic estimate with $v_{\text{out}}$ replaced with $c_{T,\text{gas}}$, $L_{\rm d}/v_{\rm esc}^{2}$, and the prediction of marchant2021 $\dot{M}_{\text{M}}$. In addition, we show also the photon tiring limit $\dot{M}_{\text{tir,L1}}$ (equation 47). Bottom right panel shows the condition for super-Eddington boost $\beta_{\text{SE,L1}}$ evaluated using $v_{\text{out}}$ or $c_{T,\text{gas}}$ (equation 48). In all panels, light and dark shading marks super-Eddington regions based on $L$ and $L_{\text{rad}}$, vertical dashed line shows the interior most point where $v_{\text{conv}}>c_{S}$, and the vertical dotted line indicates the position of the photosphere.
Models based on the porosity length formalism suggest that the maximum achievable mass loss rate is proportional to the gravitational scale height (e.g. owocki2004), and thus inversely proportional to effective gravity $g_{\rm eff}\propto\nabla\phi_{\rm R}$ in a Roche model, where $\nabla\phi_{\rm R}\to 0$ close to L1. Consequently, we assume that if an outflow occurs from a star nearly filling its Roche lobe and possessing a super-Eddington, convectively inefficient subsurface zone, it will primarily flow through the nozzle surrounding L1 at a rate $\dot{M}_{\text{out,L1}}$. The detailed steps needed for the modifications from spherical symmetry to the vicinity of L1 are described in Appendix E. In summary, we take into account the limited extent of the nozzle by multiplying $\dot{M}$ by $Q/(4\pi R_{\text{d}}^{2})\sim c_{T,\text{gas}}^{2}/v_{\text{esc}}^{2}$, and differences in potential proportional to $v_{\text{esc}}^{2}$ are multiplied by $\delta R_{\text{L}}$. These modifications give
@@
Since $\delta R_{\text{L}}$ can be made arbitrarily small by positioning the super-Eddington layer very close to the Roche lobe, this modified condition can be satisfied and the super-Eddington outflow will not be limited by photon tiring. In such a case, a super-Eddington outflow with velocity $v_{\rm out}$ and MT rate $\dot{M}_{\rm d}\sim\dot{M}_{\rm out,L1}$ forms through the L1 point. Let us note here that the outflow rate $\dot{M}_{\rm out,L1}$ does not depend on the distance from the L1 point, but only on the properties of the donor. The condition (48) is satisfied close to the L1 point because the photon-tiring limit $\dot{M}_{\text{out,L1}}$ increases as the super-Eddington layer approaches the L1 point.
quataert16 presented a different formulation of the photon tiring limit. They argued that equation (42) is valid when the critical velocity of the outflow, $v_{\text{crit}}\sim(GL_{\text{d}})^{1/5}$, exceeds the local escape velocity modified by a factor taking into account the structure of the surrounding envelope. In the vicinity of the L1 point, their condition reads $v_{\text{crit}}\gtrsim v_{\text{esc}}\delta R_{\text{L}}^{1/2}$. We show this condition in the figures that illustrate the super-Eddington MT boost, but we do not find significant differences in the outcomes when compared to equation (48).
@@ 5.2 Application to a realistic star
The higher the photon-tiring limit $\dot{M}_{\rm tir,L1}$ (equation 47), the easier it is to form an outflow through the L1 point. Therefore, we want the Eddington factor $\Gamma_{\rm Edd}$ to be as high as possible in the super-Eddington layer. The maximum Eddington factor inside stars was investigated by sanyal2015, who analyzed massive stars at LMC metallicity. They found that the local Eddington factor can exceed seven due to the hydrogen opacity bump near the photosphere in stars with effective temperatures around 5500 K (see their fig. 3). Fortunately, a similar situation arises in the $30\,\mathrm{M}_{\odot}$ low-metallicity donor studied by marchant2021.
@@
Thus, to illustrate the super-Eddington MT boost we use the same $30\,\rm{M}_{\odot}$ low-metallicity donor as in Sec. 4.4.1 but this time in a binary with an initial $4.5\,\rm{M}_{\odot}$ black hole and initial orbital period of $1000$ days. As before, we recalculated the evolution using the full MT prescription $\dot{M}_{\rm{M}}$ and the same MESA version as in marchant2021. Our first goal is to investigate how to evaluate the conditions for a super-Eddington MT boost on realistic stellar profiles. In Fig. 6, we show radial profiles of a number of relevant quantities near the outer edge of the stellar envelope at a moment, when the donor still underfills its Roche lobe. Donor radius is $R_{\rm d}=0.9963R_{\rm L}$ (vertical dotted line in Fig. 6), which means that the relative radius excess is $\delta R_{\rm d}=-0.00365$. We see that the super-Eddington layer reaches $\Gamma_{\text{tot}}\gtrsim 10$, occurs in the hydrogen recombination zone, and is accompanied by a density inversion (upper left panel). The rapid decline of $\Gamma_{\rm tot}$ near the photosphere is caused by the sharp drop in opacity (upper left panel), while the luminosity $L$ remains constant throughout the donor’s envelope (middle left panel). We find that simple estimates of the convective energy flux differ sufficiently from the results of mixing length theory (MLT) of convection implemented in MESA and are not useful. Instead, we diagnose the inefficiency of convection by comparing the $v_{\text{conv}}$ reported by MLT with the local sound speed (lower left panel), finding that convection is inefficient in a region $0.990\lesssim r/R_{\text{L}}\lesssim 0.993$. This result is insensitive to the use of $c_{S}$ or $c_{T,\text{gas}}$ for comparison purposes. The lower left panel of Fig. 6 also shows $v_{\text{crit}}$. Near the L1 point, we find that the relevant quantities for the quataert16 version of the photon tiring are $v_{\text{crit}}\approx 25\,$km s-1, $v_{\text{esc}}\approx 120\,$km s-1, and $\delta R_{\text{L}}^{1/2}\approx 0.1$ implying that the requirement is satisfied and equation (42) can be used to find $v_{\text{out}}$. Our results indicate that $v_{\text{out}}$ exceeds the local sound speed by a factor of 3–4, but remains smaller than $v_{\text{esc}}$. The middle right panel of Fig. 6 displays $\dot{M}$ near L1 corresponding to the MT flow and the photon tiring limit. The shape of $\dot{M}_{\text{tir}}$ is mostly determined by $\alpha_{\text{tir}}$ (upper right panel). We find that $\dot{M}_{\text{out,L1}}<\dot{M}_{\text{tir,L1}}$ and correspondingly $\beta_{\text{SE,L1}}<1$ in a significant range of $r$ (bottom right panel). Since there is some ambiguity about the correct value of $v_{\text{out}}$, we also show an estimate of $\dot{M}=(c_{T,\text{gas}}/v_{\text{esc}})^{2}L_{\rm d}/c_{T,\rm gas}^{2}=L_{\rm d}/v_{\text{esc}}^{2}$ and modified $\beta_{\text{SE,L1}}(c_{T,\text{gas}})$, where $v_{\text{out}}$ is replaced with $c_{T,\text{gas}}$ in equation (48). Even in this more pessimistic case, the implied MT rate does not exceed the photon tiring limit. While in this case the outflow condition is satisfied later in the evolution (only about 700 years before Roche-lobe overflow), the implied MT rates are higher leading to a higher total mass lost.
@@
Figure 7: Time evolution of donor properties illustrating super-Eddington boosted MT on an initially $30\,\rm{M}_{\odot}$ star in a binary with a $4.5\,\rm{M}_{\odot}$ black hole. Top panel shows estimates of L1 MT rates $\dot{M}_{\text{out,L1}}$ (equation 46), $\dot{M}_{\text{M}}$ (marchant2021), and a more pessimistic estimate for super-Eddington MT $\dot{M}=L_{\rm d}/v_{\text{esc}}^{2}$, as well as photon tiring limit $\dot{M}_{\text{tir,L1}}$ and wind mass loss from the donor $\dot{M}_{\text{wind}}$. Middle panel shows the evolution of relevant velocities $v_{\text{out}}$, $v_{\text{esc}}$, $v_{\text{crit}}$, and $c_{S}$. Bottom panel shows time evolution of the minimum of $\beta_{\text{SE,L1}}$, maximum of $v_{\text{conv}}/c_{S}$, and maximum of $\Gamma_{\text{tot}}=L/L_{\rm Edd}$ evaluated over the subsurface region that could drive super-Eddington MT, and the relative Roche-lobe-filling factor of the star $1-R_{\text{d}}/R_{\text{L}}$.
@@
To summarize Fig. 6, we find that at this point in evolution the donor is underfilling its Roche lobe, but the super-Eddington flux in the subsurface layer probably leads to MT through L1 with a rate of around $10^{-2}\,\rm{M}_{\odot}\,\text{yr}^{-1}$, which is significantly lower than the photon tiring limit near L1 but also significantly higher than the MT rate of about $10^{-4}\,\rm{M}_{\odot}\,\text{yr}^{-1}$ calculated using the traditional approach (marchant2021).
We now proceed to evaluate the time dependence of the super-Eddington MT boost and its importance for binary evolution by post-processing the evolution calculated using the traditional MT model without the super-Eddington boost. This serves to illustrate the possible outcomes rather than provide quantitative predictions, which would require a self-consistent evolutionary calculation. In Fig. 7, we show the time evolution of relevant quantities a few thousand years before and after Roche-lobe overflow, which occurs at about $6.973$ Myr and which we set as $t=0$. We find that the donor with the super-Eddington subsurface layer expands and the convection efficiency decreases until about t=12t=-12 kyr when $v_{\text{conv}}>c_{S}$. Shortly after, at t=11t=-11\,kyr, $\beta_{\text{SE,L1}}$ decreases below unity, and the donor should begin experiencing super-Eddington boosted MT. The implied MT rates of $10^{-3}$ to $10^{-2}$$\rm{M}_{\odot}\,\text{yr}^{-1}$ dramatically exceed the traditional optically thin MT rates or the donor’s mass loss due to stellar winds (marchant2021). Since on surface $\Gamma_{\text{tot}}\lesssim 0.1$, the Eddington boost to the optically thin MT rate discussed in Section 4.3.1 is not strong. For the final $1$ kyr before Roche-lobe overflow, the traditionally calculated MT rate quickly increases and approaches $10^{-3}$$\rm{M}_{\odot}\,\text{yr}^{-1}$. At $t=0$, the super-Eddington MT boost formally ends because the driving region moves outside the L1 point.
@@
By integrating $\dot{M}_{\text{d}}$, we find that the total mass lost from the donor due to the super-Eddington MT exceeds $10\,\rm{M}_{\odot}$. We obtain even higher numbers for the more pessimistic case when $v_{\text{out}}=c_{T,\text{gas}}$ or similar numbers even when we restrict the super-Eddington MT to the final $1$ kyr before Roche-lobe overflow when $v_{\text{crit}}>v_{\text{esc}}\delta R_{\text{L}}^{1/2}$. These numbers are clearly an artifact of not evolving the binary self-consistently. We performed preliminary MESA simulations with the super-Eddington boost included (Hadač, Cehula, Pejcha, in preparation) and found that the star responds to pre-Roche overflow mass loss similarly to standard mass transfer, leading to overall similar evolution, however, the timing and observable signatures of the mass loss are different.